the measure of angle a in triangle abc is three times the measure of angle b and the measure of angle c is half the measure of angle b find the measure of each angle
Answer by ewatrrr(24785)

3b + b + b/2 = 180
(9/2)b = 180
b = 180(2/9) = 40
a = 120
b = 40
c = 20
Sum is 180